OpenStudy (anonymous):

FAN AND MEDAL!! The sum of two opposite numbers is equal to.... (7th grade mathematics)

OpenStudy (anonymous):

answer choices?

OpenStudy (ccswims):

do you mean the sum of 2 negatives?

OpenStudy (anonymous):

I think

OpenStudy (ccswims):

a negative plus a negative = a positive outcome always

OpenStudy (anonymous):

negative plus negative isnt positive

OpenStudy (anonymous):

negative plus negative is negative

OpenStudy (anonymous):

@ccswims

OpenStudy (anonymous):

-2-4=-6

OpenStudy (anonymous):

exactly thank you @Princer_Jones

OpenStudy (anonymous):

now a negative times a negative is a positive @ccswims

OpenStudy (anonymous):

the answer is 0

OpenStudy (ccswims):

-9 - (-10) =

OpenStudy (ccswims):

1

OpenStudy (anonymous):

If a is a number, then the opposite of a is -a . hence a+(-a)=a-a=0

OpenStudy (anonymous):

@ccswims its addition not subtraction

OpenStudy (ccswims):

sorry i meant the addition sign not sub

OpenStudy (ccswims):

you know what nvm

OpenStudy (anonymous):

thats because ur distributing a negative sign!! if u werent then the answer would be negative and the way u described it u werent saying distributing

OpenStudy (ccswims):

you know what, i made a mistake and I admit that

OpenStudy (anonymous):

If a is a number, then the opposite of a is -a . hence a+(-a)=a-a=0

OpenStudy (anonymous):

A+

OpenStudy (anonymous):

@Princer_Jones